OpenStudy (anonymous):

whats the domain of square root of (x^2-81)

Parth (parthkohli):

It's a polynomial. What do you think?

OpenStudy (anonymous):

is it (-infinity, -9) and (9, infinity)

OpenStudy (anonymous):

the expretion under square root can't be negative, do x^2-81>=0, or x^2>=81 can you finish this?

OpenStudy (anonymous):

\[f(x)= \sqrt{x^2-81}\]

OpenStudy (anonymous):

answer would be 9?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Anytime you take the squareroot of a number it equals positive and negative of that integer

OpenStudy (anonymous):

(-infinity, -9] and [9, infinity)

OpenStudy (anonymous):

okay so includes the -9 and 9 thank you!!

OpenStudy (anonymous):

yes

Parth (parthkohli):

Oh, I didn't notice the "square root of."

OpenStudy (anonymous):

but @ParthKohli is right, because in the set I, it would be ok all R as domain

OpenStudy (anonymous):

thank you very much to help me understand! wasn't sure if my answer was right or wrong! thank you!

OpenStudy (anonymous):

try this one to see if u understood the concept fully \[f(x) = \sqrt{81-x^2}\]

OpenStudy (anonymous):

that would be [-9,9]

Parth (parthkohli):

Correct!

OpenStudy (anonymous):

haha thanks!

OpenStudy (anonymous):

well done!